# A Comprehensive Simulation Framework for CXL Disaggregated Memory

### v3

@@ III-B CXL Memory Expander Model
gem5 is a modular, cycle-accurate computer system simulator. Its Full-System (FS) mode executes both user-level and kernel-level instructions and models a complete system including the OS and hardware devices, thereby simulating interactions between software and hardware more realistically [27, 28, 29]. Fig. 3 illustrates the structure of an x86 platform with classic memory subsystem used in the FS simulation mode. Note that the operating flow marked with red and blue arrows will be explained in Sec. III-C. The CPU, cache, and DDR controller are placed on the coherent MemBus, while devices such as PCI host and IDE controller are mounted on the non-coherent IOBus. These two buses are connected via a bridge. In our implementation, the CXL memory expander is accessed as a PCI device attached to the IOBus.
@@
Fig. 4 shows our CXL memory expander model implemented in gem5. The model is structured from top to bottom, with a response port for communication, a memory controller for parsing read and write packets, and a memory medium for data storage. Since gem5 models external devices in a coarse-grained manner using the atomic mode (transactional behavior is represented as simple accumulated latency), we introduced the event-driven timing mode into the CXL memory expander to achieve accurate memory access behavior. The response port is connected to the IOBus, enabling it to receive request packets directed to the memory expander and send response packets to the request sources.
@@
The memory controller in the device model has two main functions. First, it parses CXL transaction packets, which include read and write configuration packets for CXL.io and memory transaction packets for CXL.mem. Second, it translates the address in the incoming packets to the actual address for accessing the backend memory medium. For the CXL.io sub-protocol, we have implemented the PCI configuration space of the expander following the PCI device programming specifications in gem5. This includes basic device information and Base Address Register (BAR) space sizes, thereby enabling the device to respond to read and write transactions from the host to the configuration space. For the CXL.mem sub-protocol, the memory controller first identifies the command fields in the packets according to the protocol specifications. Then, it calculates the actual address based on the destination address of packet and the base address stored in the BARs. The derived read/write command and its associated address are subsequently encapsulated in a new request to the memory medium for programming and retrieving data.
@@
The memory medium is designed to be technology-agnostic, which means it can be the conventional charge-based memories such as DRAM and Flash, or emerging resistance-based non-volatile memories such as RRAM and MRAM. In the case of DRAM, the memory capacity of device is declared via the BARs. During the startup of simulation on CXL-DMSim, the host machine is requested to allocate memory in the heap space of the host process based on the BAR space capacity specified in the Python script. Afterwards, upon receiving read and write requests from the memory controller, the memory medium accesses the data at the corresponding address and generates responses if needed. For the memory medium, we currently offer two DRAM models. The first is a coarse-grained model that offers a faster simulation at the cost of some accuracy. The second is the native DRAM model from gem5, which provides higher simulation accuracy but operates at a slower speed. Additionally, thanks to the modular design of gem5, other existing memory models, such as the NVM model and DRAMSim model, can be easily integrated into our expander model to accommodate the research needs of various users.
Fig. 5 shows the modeled key contributions of latency for a CXL.mem access request from CPU and its response from the CXL memory expander in CXL-DMSim. To capture the timing behavior of CXL memory expanders, our CXL memory expander model provides two parameters to simulate the processing latency of the CXL.mem sub-protocol and the access latency of the memory medium: device_proto_proc_lat and medium_access_lat. Note that the other two key latency parameters bridge_1at and host_proto_proc_lat associated with the bridge in the figure will be explained later in Sec. III-C.
@@ III-C CXL Protocol Support
The CXL.mem sub-protocol defines two types of communication endpoints. One is called master, typically a local agent in the host processor. The other is subordinate such as a memory controller in the CXL memory expander. This protocol specifies two types of packets for communication between the master and subordinate agents: those from the master to the subordinate are called Master-to-Subordinate (M2S) packets, and those in the opposite direction are called Subordinate-to-Master (S2M) packets [16].
CXL-DMSim extends the packet-based point-to-point communication mechanism in gem5 to support the simulation of the CXL.mem sub-protocol. First, when the CPU launches a ReadReq/WriteReq gem5 packet, its command field is updated to include a read/write transaction command used by the CXL.mem sub-protocol. Subsequently, the bridge connecting the memory bus and I/O bus intercepts the packet if destined to the CXL memory expander. Fig. 6 shows that the bridge contains two pairs of FIFO queues, one pair for buffering upstream gem5 packets and the other pair for buffering downstream CXL.mem packets. The bridge checks the address and captures the gem5 packet in the upstream Req FIFO, then transforms it into a M2SReq/M2SRwD packet. Thereafter, the CXL.mem packet is buffered in the downstream Req FIFO and then sent out to the CXL memory expander. Similarly, when a response packet (S2MDRS or S2MNDR) from the CXL memory expander arrives at the bridge, it is buffered in the Resp FIFO and then converted into a ReadResp/WriteResp packet. The gem5 packet is buffered in the upstream Resp FIFO before being transmitted to the memory bus.
Note that the bridge module includes four configuration parameters (see Fig. 5): bridge_lat, host_proto_proc_lat, req_fifo_depth, and resp_fifo_depth. The bridge_lat represents the inherent latency through the bridge module, which is already present in the original gem5 code. The host_proto_proc_lat represents the processing latency for CXL.mem sub-protocol packets. The req_fifo_depth and resp_fifo_depth determine the depths of the FIFO queues used for buffering CXL.mem request and response packets, respectively. These queue depths impact the CXL memory expander’s capability to handle concurrent requests.
@@ IV-A1 Hardware Testbed
On our hardware testbed, we measured two types of CXL memory devices: a real CXL FPGA memory device (denoted as CXL-FPGA) and a real CXL ASIC memory device (denoted as CXL-ASIC); the details are listed in Table II. Both of these two devices are connected to Socket1, as illustrated in Fig. 9. The CXL-FPGA device is an Intel Agilex I-Series FPGA Development Kit with an integrated hard CXL IP [30]. It is equipped with 16GB 3200MT/s DDR4 memory, connected to the CPU via an CXL 1.1 interface. The CXL-FPGA device is recognized as a separate CPU-less NUMA node (denoted as NUMA4); it can be used and managed using existing NUMA software infrastructure. The in-house CXL-ASIC device has two CXL memory controllers, each of which is composed of a CXL protocol controller plus a DDR5 controller driving two DIMMs. As each DIMM is populated with a 32GB DDR5 4800MT/s memory module, the CXL-ASIC device owns a total memory capacity of 128GB. It appears in the OS as two separate NUMA nodes without CPUs (denoted as NUMA2 and NUMA3 respectively).
We also measured the DDR5 memory on the remote NUMA node (denoted as DDR-R on NUMA0) as an emulated CXL memory device, as done in many of the prior works [21, 22, 23]. For better comparing the performance of the above-mentioned three types of memory, we also measured the local DDR5 memory (denoted as DDR-L) on NUMA1 as a baseline.
@@
| Config. Parameter | CXL-DMSimF | CXL-DMSimA |
| CXL memory size | 16GB | 64GB |
| bridge_lat | 50ns | 50ns |
| host_proto_proc_lat | 14ns | 14ns |
| device_proto_proc_lat | 60ns | 15ns |
| medium_access_lat | 50ns | 50ns |
| req_fifo_depth | 48 | 52 |
| resp_fifo_depth | 48 | 52 |
@@ IV-A2 CXL-DMSim Simulator
CXL-DMSim is implemented based on gem5 v23.1 in the full-system mode. The CPU type is X86O3CPU with 32GB 4400MT/s single DDR5 memory (denoted as CXL-DMSimL). The CXL memory device driver was developed based on the Linux v5.4.49 kernel. CXL-DMSim simulates a CXL FPGA memory device (denoted as CXL-DMSimF) with 16GB and a CXL ASIC memory device (denoted as CXL-DMSimA) with 64GB, with configurations listed in Table III. To simulate different CXL memory devices from different vendors, we introduce six configurable parameters and their typical ranges in our simulator, according to the latency breakdown of memory access paths provided in the CXL spec. [16] and relevant literature [31, 3, 32, 33, 34, 35, 9]. For our CXL FPGA and ASIC devices, the values of these parameters in the Table III have been meticulously calibrated and fine-tuned using end-to-end latency measurements on our hardware testbed. For those who are interested in using our simulator, different values can be configured to match the performance of their own CXL memory devices.
@@ IV-B Usability and Fidelity of CXL-DMSim
There exist many popular benchmarks to characterize memory performance in industry, such as Intel’s MLC [36], LMbench [37], STREAM [38]. Considering the ease of use and compatibility on both hardware and simulator platforms, we selected LMbench, STREAM, and Redis-YCSB [39, 40] to evaluate the latency, bandwidth, and real-world application performance of DDR-L, DDR-R, CXL-FPGA, CXL-ASIC and CXL-DMSimF, and CXL-DMSimA.
@@ IV-B1 LMbench Test for Memory Latency
To measure the access latency of CXL HDM, we utilized a memory latency benchmark lat_mem_rd in LMbench [37]. It can be used to measure random read latency of different memory layers, covering L1, L2, and L3 caches as well as DDR-L, DDR-R, and CXL memories. The random read pattern typically reflects the system’s real latency [41, 42]. Hardware prefetch mechanisms are all disabled in BIOS settings to provide more accurate measurements of memory latencies. Note that we only present the read latency as the CXL memory is symmetric in read and write performance. The benchmark involves two key parameters: array size and stride. To ensure that the entire memory hierarchy is accurately measured, avoiding the disturbance of caching on the results meanwhile properly representing the data access patterns in real-world applications, the array size should be set to at least four times as large as LLC and no more than 80% of physical memory size [43]. Given the above considerations, we set the array size to 1024MB and the stride to 64 bytes in our experiments.
Fig. 10 shows the measurement results using LMbench. The curves contain a series of plateaus, each of which represents a level in the memory hierarchy. The right-most plateau represents the random read latency of DDR-L (), DDR-R (), CXL-FPGA (), CXL-ASIC (), CXL-DMSimF (), and CXL-DMSimA (). The latency of CXL-FPGA is approximately 2.88 times higher than that of DDR-L, while the latency of CXL-ASIC is about 2.18 times. The two types of CXL memory devices exhibit much higher latencies, because the access path of CXL memory is much longer than that of DDR-L. The latency of CXL-FPGA is higher than CXL-ASIC, because FPGA-based CXL memory fails in fully utilizing DRAM chip performance due to FPGA’s lower operating frequency compared to ASICs [44]. It is worth noting that the latencies between 192MB and 384MB for CXL-DMSim are consistently higher than those measured on real hardware. This discrepancy is primarily attributed to: 1) poor CPU model in gem5 lagging behind sophisticated real-world CPUs, 2) smaller LLC size (96MB due to restricted value selection) in the gem5 setting than that (97.5MB) on the real hardware platform. When using remote NUMA to emulate CXL memory, the read latency is relatively lower compared to real CXL memory devices, being about 1.53 times higher than that of DDR-L. Clearly, the experimental results show that CXL-DMSim can accurately simulate the access latency of real CXL memory devices, while the remote NUMA emulation cannot reflect the real latency of the devices.
@@ IV-B2 STREAM Test for Memory Bandwidth
To measure the sustainable memory bandwidth of our prototyped and simulated systems, we used STREAM [38] as a benchmark. As a well-known benchmark, STREAM provides four distinct tests, i.e., Copy, Scale, Add, and Triad with different memory access and calculation patterns.
@@
To ensure that STREAM measures realistic bandwidth between memory and processor without interference from caches, the data size has to be at least 8 times of the LLC size; thus we set the data size to 1024MB in our experiments. Fig. 11 shows the measurement results of the six different memory devices. Note that all the values are normalized to the bandwidth of DDR-L for the purpose of better comparison. It can be seen that the bandwidths derived from CXL-DMSim closely align with those obtained from the CXL hardware prototype; the modeling error rate is about 6% on average. The CXL-FPGA memories achieve approximately 45%-69% of the DDR-L bandwidth. Comparatively, the CXL-ASIC memories exhibit a better performance in bandwidth, achieving about 82%-83% of the DDR-L bandwidth. One can also see that the bandwidth of DDR-R is 68%-74% of the DDR-L bandwidth. This indicates that the bandwidth of DDR-R memory lies between CXL-ASIC and CXL-FPGA memories.
The reason why the measured bandwidth of the emulated CXL memory device using remote DDR5 memory (i.e., DDR-R) outperforms the real CXL FPGA memory device can be explained as follows. While CXL 1.1 is based on PCIe 5.0 x16 achieving a maximum theoretical bi-directional bandwidth of 128GB/s, the on-board DDR4 memory on our FPGA platform is limited to a maximum bandwidth of 25.6GB/s. The comparatively low bandwidth of DDR4 memory actually poses the bottleneck of the real CXL memory device. When it comes to the CXL-ASIC device, it is equipped with two channels of DDR5 memory with 76.8GB/s, which significantly enhances the CXL link’s actual bandwidth in comparison to the DDR4 memory on the FPGA board. Furthermore, the ASIC’s sophisticated architecture and elevated operating frequency also contribute to its higher bandwidth.
IV-B3 Mess Test With Varying Rd/Wr Patterns
We also conducted tests on our platform using the newly-proposed Mess benchmark [26]. Fig. 12(a) illustrates that DDR-L memory exhibits a marginal variation in bandwidth across different Rd/Wr patterns, whereas Fig. 12(b) demonstrates that CXL-ASIC memory is much more sensitive to Rd/Wr patterns (approximately as marked with red two-way arrows). This discrepancy arises from the fact that the DDR protocol utilizes a bi-directional parallel bus for both reads and writes in a time-multiplexed manner, while the CXL protocol features SerDes TX/RX expressways for separated read and write traffics.
In addition, one can observe that the bandwidth of DDR-L maximizes at the 100% read pattern, and it gradually decreases until reaching at Rd:Wr=50%:50%. In contrast, CXL-ASIC exhibits the best performance at Rd:Wr=74%:26%, instead of 50%:50% as claimed in the Mess work [26]. we argue that this phenomenon is caused by the compromised CXL memory controller design, where the CXL protocol has to be translated to DDR for accessing DRAM storage arrays. This CXL+DDR design not only prolongs memory access paths leading to increased access latency, but also compromises the bandwidth utilization of CXL link and power consumption. Hence, we believe it is necessary to eliminate DDR in the CXL memory controller by directly exposing backend DRAM arrays to the CXL protocol, similar to what NVMe has done to SATA SSDs.
IV-B4 Real-World Application Test

### v6

@@ III-B CXL Memory Expander Model
gem5 is a modular, cycle-level computer system simulator. Its full-system (FS) mode executes both user-level and kernel-level instructions and models a complete system including the OS and hardware devices, thereby simulating interactions between software and hardware more realistically [23, 24]. Fig. 3 illustrates the structure of an x86 platform with the classic memory subsystem used in the FS simulation mode. Note that the operating flow marked with red and blue arrows will be explained in Sec. III-C. The CPU, cache, and DDR controller are placed on the coherent MemBus, while devices such as PCI host and IDE controller are mounted on the non-coherent IOBus. These two buses are connected via a bridge. In our implementation, the CXL memory expander is accessed as a PCI device attached to the IOBus.
@@
Fig. 4 shows our CXL memory expander model in CXL-DMSim. The model is organized into four components from top to bottom: 1) a response port for communication, 2) a CXL controller for parsing read/write packets, 3) a flexible interface layer for request forwarding, and 4) a memory module including memory controllers and memory medium for data storage. Since native gem5 models external devices in a coarse-grained manner using the atomic mode (transactional behavior is represented as simple accumulated latency), this approach loses critical timing information essential for memory modeling. To achieve accurate memory access behavior, we introduce an event-driven timing mode into the CXL memory expander. The response port is connected to the IOBus, enabling it to receive request packets directed to the memory expander and send response packets to the requesters.
@@
The CXL controller in the device model incorporates Req FIFO and Rsp FIFO buffers to manage incoming request packets from the port and response packets from the backend memory medium, respectively (note that the definitions of M2SReq, M2RRwD, S2MDRS, and S2MNDR packets shown in Fig. 4 will be explained in Sec. III-C). The CXL controller performs two critical operations on buffered packets. First, it parses CXL transaction packets, which include read and write configuration packets for CXL.io and memory transaction packets for CXL.mem. Second, it translates the address in the incoming packets to the actual address for accessing the backend memory controller. For the CXL.io sub-protocol, we have implemented the PCI configuration space of the expander. This includes basic device information and base address registers (BARs), thereby enabling the device to respond to read and write transactions from the host to the configuration space. For the CXL.mem sub-protocol, the CXL controller first identifies the command fields in the packets according to the protocol specifications. Then, it calculates the actual address based on the destination address of the packet and the base address stored in the BARs. The derived read/write commands and their associated addresses are encapsulated into new requests, which are forwarded to the memory controller through the interface layer to execute load/store operations. This flexible interface layer decouples the frontend CXL controller from the backend memory module, enabling convenient customization of CXL memory characteristics such as latency, bandwidth, and volatility through memory medium replacement and adjustment of memory channel count.
@@
The memory module is designed to be technology-agnostic, which means it can be the conventional charge-based memories such as DRAM and Flash, or emerging resistance-based non-volatile memories such as RRAM and MRAM. The memory capacity of the device is declared to the host via the BARs during system startup. When the memory controller’s transaction queue receives read and write requests from the CXL controller, it schedules the requests to access data at the corresponding addresses and generates responses if needed. Thanks to our well-designed modular structure, existing memory models from gem5, including native DRAM models, DRAMSim models, and NVM models, can be directly integrated into our expander’s backend. Additionally, we provide a coarse-grained DRAM model that offers faster simulation speed at the cost of reduced accuracy.
To capture the timing behavior of CXL memory expanders, our device model provides three key parameters: dev_proto_lat regulates the processing latency of the CXL.mem sub-protocol, while dev_req_fifo_depth and dev_rsp_fifo_depth determine the CXL controller’s capacity to handle concurrent packets via FIFO depth configuration. In the timing mode, the overall memory access latency and expander bandwidth are functions of packet queuing status, memory controller scheduling policies, and the temporal adjacency of memory requests.
@@ III-C CXL Protocol Support
The CXL.mem sub-protocol defines two types of communication endpoints. One is called master, typically a local agent in the host processor. The other is subordinate such as a memory controller in the CXL memory expander. This protocol specifies two types of packets for communication between the master and subordinate agents: those from the master to the subordinate are called master-to-subordinate (M2S) packets, and those in the opposite direction are called subordinate-to-master (S2M) packets [11]. CXL-DMSim extends the packet-based point-to-point communication mechanism in gem5 to simulate the CXL.mem sub-protocol. First, when the CPU issues a ReadReq/WriteReq gem5 packet, its command field is updated to include a read/write transaction command used by the CXL.mem. Subsequently, the bridge connecting the memory bus and I/O bus intercepts the packet if destined to the CXL memory expander. Fig. 5 shows that the bridge contains two pairs of FIFO queues, one pair for buffering upstream gem5 packets and the other for buffering downstream CXL.mem packets. The bridge checks the address and captures the gem5 packet in the upstream Req FIFO, then transforms it into a M2SReq/M2SRwD packet. Thereafter, the CXL.mem packet is buffered in the downstream Req FIFO and then sent out to the CXL memory expander. Similarly, when a response packet (S2MDRS/S2MNDR) from the CXL memory expander arrives at the bridge, it is buffered in the Rsp FIFO and then converted into a ReadRsp/WriteRsp packet. The gem5 packet is buffered in the upstream Rsp FIFO before being transmitted to the memory bus.
We also implement a CXL switch model to support the single-level switching feature introduced in CXL2.0 protocol. With CXL switches, a larger disaggregated memory system can be built as illustrated Fig. 3. In single-host configurations, a CXL switch contains a single upstream port (USP) and several downstream ports (DSPs). Each port corresponds to a virtual PCI-to-PCI bridge (vPPB), and these vPPBs are interconnected through a virtual PCI bus for packet forwarding. We model the CXL switch by extending the Bridge component. When the USP receives an upstream packet, its vPPB inspects the packet’s address and forwards the packet to the corresponding DSP. Similarly, a packet routed from an endpoint device to the host goes through the switch from the device’s associated DSP to the USP in a reverse direction.
Note that the bridge module includes four configuration parameters: The bridge_lat represents the inherent latency through the bridge module, which is already present in the original gem5 code. The host_proto_lat represents the processing latency for CXL.mem sub-protocol packets. The link_req_fifo_depth and link_rsp_fifo_depth determine the depths of the FIFO queues used for buffering CXL.mem request and response packets, respectively. These queue depths impact the throughput of the CXL link. Similarly, the CXL switch includes two configurable parameters: switch_lat represents the packet forwarding latency through the switch, and switch_buffer_size represents the size of the switch buffer.
@@ IV-A1 Hardware Testbed
CXL-FPGA and CXL-ASIC devices.
@@
On our hardware testbed, we measured two types of CXL memory devices: a real CXL FPGA memory device (denoted as CXL-FPGA) and a real CXL ASIC memory device (denoted as CXL-ASIC); the details are listed in Table II. Both of these two devices are connected to Socket1, as illustrated in Fig. 8. The CXL-FPGA device is an Intel Agilex I-Series FPGA Development Kit with an integrated hard CXL IP [25]. It is equipped with 16GB 3200MT/s DDR4 memory, connected to the CPU via a CXL 1.1 interface. The CXL-FPGA device is recognized as a separate CPU-less NUMA node (denoted as NUMA4); it can be used and managed using existing NUMA software infrastructure. The in-house CXL-ASIC device has two CXL memory controllers, each of which is composed of a CXL protocol controller plus a DDR5 controller driving two DIMMs. As each DIMM is populated with a 32GB DDR5 4800MT/s memory module, the CXL-ASIC device owns a total memory capacity of 128GB. It appears in the OS as two separate NUMA nodes without CPUs (denoted as NUMA2 and NUMA3 respectively).
We also measured the DDR5 memory on the remote NUMA node (denoted as DDR-R on NUMA0) as an emulated CXL memory device, as done in many of the prior works [18, 19]. For better comparing the performance of the above-mentioned three types of memory, we also measured the local DDR5 memory (denoted as DDR-L) on NUMA1 as a baseline.
@@ IV-A2 CXL-DMSim Simulator
CXL-DMSim is implemented based on gem5 v23.1 in the full-system mode. The CPU type is X86O3CPU with 32GB 4400MT/s single DDR5 memory (denoted as CXL-DMSimL). The CXL memory device driver was developed based on the Linux v5.4.49 kernel. CXL-DMSim simulates a CXL FPGA memory device (denoted as CXL-DMSimF) with 16GB, a CXL ASIC memory device (denoted as CXL-DMSimA) with 64GB, and a CXL ASIC memory device interconnected via a switch (denoted as CXL-DMSimS), with configurations listed in Table III. To simulate different CXL memory devices from different vendors, we introduce nine configurable parameters and their typical ranges in our simulator, according to the latency breakdown of memory access paths provided in the CXL spec. [11] and relevant literature [26, 2, 27, 20, 28]. For our CXL FPGA and ASIC devices, the values of these parameters in Table III have been meticulously calibrated and fine-tuned using end-to-end latency measurements on our hardware testbed. For the switch-attached CXL ASIC device, we calibrated the switch’s parameters to match the performance measurements reported by XConn [28]. For those who are interested in using our simulator, different values can be configured to match the performance of their own CXL memory devices.
| Config. Parameter | CXL-DMSimF | CXL-DMSimA | CXL-DMSimS |
| CXL memory size | 16GB | 64GB | 64GB |
| Type of backend DRAM | DDR4 3200 | DDR5 4400 | DDR5 4400 |
| bridge_lat | 50ns | 50ns | 50ns |
| host_proto_lat | 12ns | 12ns | 12ns |
| dev_proto_lat | 60ns | 15ns | 15ns |
| link_req_fifo_depth | 128 | 128 | 128 |
| link_rsp_fifo_depth | 128 | 128 | 128 |
| dev_rsp_fifo_depth | 36 | 48 | 48 |
| dev_rsp_fifo_depth | 36 | 48 | 48 |
| switch_lat | N/A | N/A | 100ns |
| switch_buffer_size | N/A | N/A | 64 |
@@ IV-B Usability and Fidelity of CXL-DMSim
There exist many popular benchmarks to characterize memory performance in industry, such as Intel’s MLC [29], LMbench [30], STREAM [31]. Considering ease of use and compatibility on both hardware and simulator platforms, we selected LMbench, STREAM, and Redis-YCSB [32, 33] to evaluate the latency, bandwidth, and real-world application performance of DDR-L, DDR-R, CXL-FPGA, CXL-ASIC, and CXL-DMSimF, CXL-DMSimA, and CXL-DMSimS.
@@ IV-B1 LMbench Test for Memory Latency
To measure the access latency of CXL HDM, we utilized a memory latency benchmark lat_mem_rd in LMbench [30]. It can be used to measure the random read latency of different memory layers, covering L1, L2 and L3 caches, as well as DDR-L, DDR-R and CXL memories. The random read pattern typically reflects the system’s real latency [13]. Hardware prefetch mechanisms are all disabled in BIOS settings to provide more accurate measurements of memory latencies. Note that we only present the read latency as the CXL memory is symmetric in read and write performance. The benchmark involves two key parameters: array size and stride. To ensure that the entire memory hierarchy is accurately measured, avoiding the disturbance of caching on the results meanwhile properly representing the data access patterns in real-world applications, the array size should be set to at least four times as large as LLC and no more than 80% of physical memory size [30]. Given the above considerations, we set the array size to and the stride to 64 bytes in our experiments.
Fig. 9 presents the measurement results using LMbench. The curves contain a series of plateaus, each of which represents a level in the memory hierarchy. The right-most plateau represents the random read latency of DDR-L (), DDR-R (), CXL-ASIC (), CXL-FPGA (), CXL-DMSimL (), CXL-DMSimA (), CXL-DMSimF (), and CXL-DMSimS (). The latency of CXL-ASIC is approximately 2.18 times higher than that of DDR-L, while the latency of CXL-FPGA is about 2.88 times. The two types of CXL memory devices exhibit much higher latencies because their access path is much longer than that of DDR-L. The latency of CXL-FPGA is higher than CXL-ASIC, because FPGA-based CXL memory fails in fully utilizing DRAM chip performance due to FPGA’s lower operating frequency compared to ASICs. For the switch-attached CXL-ASIC, the measured latency is 3.75 times larger than that of DDR-L; the additional delay compared to direct-attached CXL-ASIC is introduced by the CXL switch.
Notably, the latencies between and for CXL-DMSim are consistently higher than those measured on real hardware. This discrepancy is primarily due to gem5’s LLC model not capturing the full complexity of real LLC behavior. Specifically, when the pointer-chase working set goes up above the LLC size () but remains below four times (), the observed latencies reflect a mixed effect of LLC hits and misses. Because gem5’s LLC model per se exhibits a high error rate (recent work [34] reports up to 56.4% modeling error versus just 6.2% for its DRAM model), CXL-DMSim poorly reproduces the intermediate-range latency measured in real hardware. A secondary influential factor is that the simulated LLC size is smaller than that on our testbed due to gem5’s LLC configuration constraints; a higher setting (e.g., ) leads to a smaller gap.
When using remote NUMA to emulate CXL memory, the read latency is relatively lower compared to real CXL memory devices, being about 1.53 times higher than that of DDR-L. Clearly, the experimental results show that CXL-DMSim can accurately simulate the access latency of real CXL memory devices, while the remote NUMA emulation cannot reflect the real latency of the devices.
@@ IV-B2 STREAM Test for Memory Bandwidth
To measure the sustainable memory bandwidth of our prototyped and simulated systems, we used STREAM [31] as a benchmark, which provides four tests, i.e. Copy, Scale, Add and Triad with different memory access and computing patterns. To ensure that STREAM measures realistic bandwidth between memory and processor without interference from caches, the data size has to be at least four times the size of LLC; thus we set the data size to in our experiments.
@@
Fig. 10 shows the measurement results of the seven different memory devices. Note that all the values are normalized to the bandwidth of DDR-L for the purpose of better comparison. It can be seen that the bandwidths derived from CXL-DMSim closely align with those obtained from the CXL hardware prototype; the modeling error rate is about 4.3% on average. The CXL-FPGA memory achieves approximately 45%-69% of the DDR-L bandwidth. Comparatively, the CXL-ASIC memory exhibits better performance in bandwidth, achieving about 82%-83% of the DDR-L bandwidth. One can also see that the bandwidth of DDR-R is 68%-74% of the DDR-L bandwidth. This indicates that the bandwidth of DDR-R memory lies between CXL-ASIC and CXL-FPGA memories. The simulated switch-attached CXL-ASIC (CXL-DMSimS) shows similar bandwidth to the direct-attached CXL-ASIC.
In a nutshell, the relationship of the measured bandwidth is CXL-LCXL-ASICDDR-R albeit the DDR5 bandwidth is the same in all three cases. This can be explained as follows. In the DDR-R case, UPI uses a directory-based home snoop coherency protocol [35]. A cache-line request from the local core to remote memory (DDR‑R) is first checked by the local caching and home agent (CHA) to determine whether any copies exist within the NUMA node (by consulting the snoop filter and LLC). In case none is found, it is forwarded to the remote (home) node’s CHA for cross‑node cache‑coherence checking. The memory directory tracking remote cache-line state on DRAM chips and its on-die cache (HitME) [36] are also involved to maintain coherence across NUMA nodes [37]. Under STREAM tests with mixed read/write operations, the frequent cross-UPI coherence checks, especially on writes, become a bottleneck limiting DDR-R bandwidth [38]. In contrast, the CXL-ASIC memory is physically located at the same socket as the DDR-L memory, which does not involve cross-UPI cache coherency. The reason why the bandwidth of CXL-ASIC is smaller than DDR-L is owing to the overhead of packing and unpacking CXL protocol packets as well as the current immature design of the CXL memory controller [26]. Finally, CXL-FPGA exhibits the lowest bandwidth due to the memory-bandwidth limitations of its on-device DDR4 memory and the inherently lower operating frequency of the FPGA.
IV-B3 Real-World Application Test